Which expression has the same value as 36*19
a) (30+19)*(6+19)
b) (30*10)+(6*9)
c) (30*19)+(6*19)
d) (30+10)*(6+9)

Answer by jim_thompson5910(35256)

You can express 36 as 30+6, so


36*19 = (30+6)*19 = (30*19)+(6*19)


So the answer is choice C)